# TOB-SVD: Total-Order Broadcast with Single-Vote Decisions in the Sleepy Model

### v2

@@ 7 Conclusions
Our main contribution in this work is the introduction of TOB-SVD, a novel Total-Order Broadcast protocol that supports dynamic participation and operates within (a variant of) the sleepy model. Our protocol is resilient against up to 1/2 adversarial participants and, compared to protocols with the same level of adversarial resilience, it improves both expected latency and transaction expected latency, and reduces the number of voting phases per new block to only one in the best case and to only two in the expected case. This makes our protocol significantly more practical for use in large-scale networks such as permissionless blockchain systems. Additionally, our protocol is arguably much simpler. However, the trade-off is that it requires a stabilization time of $2\Delta$ in addition to the stabilization time required by other protocols to retrieve information upon waking up. Therefore, in cases where a lot of information needs to be retrieved, such as in blockchain systems, our protocol requires only a slightly longer assumption on the overall stabilization time.

### v3

@@ 7 Conclusions
We introduce TOB-SVD, a novel Total-Order Broadcast protocol that supports dynamic participation and operates within (a variant of) the sleepy model. TOB-SVD tolerates up to 1/2 Byzantine validators and improves both expected latency and transaction latency compared to existing protocols with the same resilience, requiring only one voting phase in the best case and two in expectation. This makes it significantly more practical for large-scale networks like permissionless blockchains. The protocol is also notably simpler, at the cost of an added $2\Delta$ stabilization time on top of the time required by other protocols to retrieve information upon waking. Therefore, in scenarios where a large amount of information must be retrieved, such as in blockchain systems, our protocol requires only a slightly stronger assumption on the overall stabilization time.